OpenStudy (anonymous):

Find the product of (x + 7)(x − 7).

OpenStudy (ash2326):

Use this relation \[(a+b)(a-b)=a^2-b^2\] here a=x and b=7 Can you do it now?

OpenStudy (anonymous):

Yep Thanxsssssssssssss